OpenStudy (brooke10239):

Solve the proportion m over 63 = 2 over 9 (m/63 = 2/9)

OpenStudy (brooke10239):

@purple_pink

OpenStudy (purple_pink):

hmm

OpenStudy (purple_pink):

with that link i put out there for you, m = 2 63 9 Cross multiply: m * 9 = 63 * 2 Simplifying m * 9 = 63 * 2 Reorder the terms for easier multiplication: 9m = 63 * 2 Multiply 63 * 2 9m = 126 Solving 9m = 126 Move all terms containing m to the left, all other terms to the right. Divide each side by '9'. m = 14 Simplifying m = 14

OpenStudy (purple_pink):

i think you are afk

OpenStudy (purple_pink):

hope to get a medal